OpenStudy (anonymous):

Brian has no more than 16 coins. They comprise of both nickels and dimes. There are at least four times as many nickels as there are dimes. Which of the following is a system of inequalities that describes this situation?

OpenStudy (anonymous):

Help me?

OpenStudy (anonymous):

@radar ?

OpenStudy (anonymous):

@radar Did you fall asleep on the keyboard?

OpenStudy (radar):

I am reading it and trying to make some sense of it: One thing that it is stating is the sum of the number of nickels and dimes is less than 16. So n + d < 16 , where n is number of nickels and d is the number of dimes. But there is more that must be considered. But, are you with me so far?

OpenStudy (anonymous):

Yes.

OpenStudy (radar):

No lol, I just had a nap.

OpenStudy (anonymous):

Oh okay lol

OpenStudy (radar):

Oh they are also stating that 4n=d four times the number of nickels as their are dimes. Now let see how this little bit of information is used.

OpenStudy (anonymous):

Okie dokie.

OpenStudy (radar):

Well I guess it could look like this 4n + d < 16. Let us assume an example if he had 3 dimes and 12 nickels he would have less then 15. That is an example of a scenario that falls within the inequality. However if he had 4 dimes and 16 nickels that could not be a possibility as he has 20 coins and that is a no no, (more than 16)

OpenStudy (anonymous):

So the answer would be?

OpenStudy (radar):

change the 15 to 16. (typo) the answer would be 4n + d < 16

OpenStudy (radar):

What are the options?

OpenStudy (radar):

You say following systems, what are the systems?

OpenStudy (anonymous):

\[A: D+N \le16 4D \le N or B: D+N \ge16 4D \ge Nor C:D+N \le16 4D \le-Nor D:D+N \le16 $D geN\]n

OpenStudy (radar):

It is not B, it is not C It is A as the statement says there are at least 4 times the number of nickels than dimes. I misread it, I thought it said there were 4 times the number of nickels than dimes Actually I would think the best answer would be: D + N < 16 and4N <=D

OpenStudy (radar):

Was there a choice D?

OpenStudy (anonymous):

\[D: D+N \le 16 and 4D \ge N\]

OpenStudy (anonymous):

Hello?

OpenStudy (radar):

No question that the \[D + N \le16\] part is correct . But I am getting twisted on the "there are at least four times as many nickels as dimes" I guess they are saying there could be as many as 4, 3, 2, 1 times but not more than four......so \[D \le4N\] But do you have an option that combines those two?

OpenStudy (anonymous):

No.

OpenStudy (radar):

I think A is saying the same thing in a different manner.

OpenStudy (anonymous):

So it's A?

OpenStudy (radar):

Sorry I can not be more definitive, but I like A

OpenStudy (radar):

I couldn't see all of C. What is choice C.? I think I see a minus sign.

OpenStudy (anonymous):

I turned it in already.

OpenStudy (radar):

\[-D \ge-4n\] is the same as the way I expressed Maybe that is what option C had, I couldn't read option C.

OpenStudy (anonymous):

Okay.

OpenStudy (radar):

Was A correct?

OpenStudy (radar):

Or which one did you select?

OpenStudy (anonymous):

I don't remember lol